# To Buy an Electric Vehicle or Not? A Bayesian Analysis of Consumer Intent in the United States

### v1

@@ 3 Data
The study draws on survey data from three opinion polls carried out by the Pew Research Center over a three-year span. The first survey was conducted during April 20-21, 2021 on a sample of 13,749 adults, the second survey took place during May 2-8, 2022, on a sample of 10,282 adults, and the third survey was administered during May 30-June 4, 2023 on a sample of 10,329 adults. After removing all missing observations and cleaning the data, we have a sample of 11,052, 7,173, and 7,201 observations available for analysis. Table 1 shows a descriptive summary of the independent and dependent variables. Specifically, it presents the mean and standard deviation of the continuous variables, and count and percentage for the categorical variables.
The above survey data offer at least three distinct advantages over existing studies that use survey data to gauge EV purchase intent. First, our survey data is nationally representative and span multiple years. As a result, the estimates we generate are more likely to be robust to cross-sectional heterogeneity. All survey data and additional information on the surveys used in this paper are available for download from the Pew Research Center dataset portal (and the corresponding author’s webpage). Second, an appealing feature of our data sets is the availability of ordinal responses to likelihood of purchasing EV. This enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as (Buhmann and Criado, 2023). Third, we note that the first and second surveys correspond to the Covid-19 period, while the third survey was conducted in the post-pandemic phase. Therefore, if there was an increase in interest for sustainable products like electric vehicles during the pandemic, this should be evident in the data summary, with a higher percentage of respondents expressing a preference for buying electric vehicles during the pandemic phase.
@@
EV Info $\circ$ Read or heard a lot 4170 37.73 ..absent... . ..absent... . ..absent... . ..absent... .
$\circ$ Little or nothing at all 6882 62.27 ..absent... . ..absent... . ..absent... . ..absent... .
Env Better $\circ$ EVs are environmentally ‘better’ than gas-powered vehicles 7944 71.88 ..absent... . ..absent... . ..absent... . ..absent... .
$\circ$ About the same or worse 3108 28.12 ..absent... . ..absent... . ..absent... . ..absent... .
EV Infra $\circ$ US will build charging stations and infrastructure: extremely or very confident ..absent... . ..absent... . ..absent... . ..absent... . 1263 17.54
$\circ$ Somewhat confident, not too confident, or not at all confident ..absent... . ..absent... . ..absent... . ..absent... . 5938 82.46
@@
Married Married or living with a partner ..absent... . ..absent... . 5692 79.35 5640 78.32
@@
We begin by discussing the independent variables that have a direct influence on purchasing an EV–the question of primary interest in this article. Two such variables, ‘EV Info’ and ‘Env Better’, are specific to the April 2021 survey. The variable ‘EV Info’ measures how much a respondent has read or heard about EVs. It is coded as 1 if the respondent has read or heard a lot about EVs, and 0 if they have read or heard little or nothing. As shown in Table 1, about 38% of respondents have read or heard a lot about EVs. The variable ‘Env Better’ reflects perception about the environmental impact of EVs as compared to gas-powered vehicles. This indicator variable takes the value 1 if the respondent believes that EVs are better for the environment than gas-powered vehicles, and 0 otherwise (worse for the environment or about the same). According to Table 1, approximately 72% of respondents view EVs as being environmentally superior to gas-powered vehicles. Another noteworthy variable is ‘EV Infra’, which is only available in the May-June 2023 survey. This indicator variable is coded as 1 if the respondent is extremely or very confident that the US will build the necessary charging stations and infrastructure to support large number of EVs on the roads, and 0 otherwise (somewhat confident, not too confident, or not at all confident). As seen in Table 1, about 82% of respondents express a negative perception towards the government for creating the infrastructure needed to support the adoption of EVs.
@@
Two key variables of interest to our study, which are available in all the surveys, are ‘EV Owner’ and ‘GACC’. The ‘EV Owner’ variable indicates whether the respondent currently owns an electric or hybrid vehicle and takes the value 1 if true and 0 otherwise. As shown in Table 1, the percentages of respondents who are current owners of an electric or hybrid vehicle are 9.40%, 10.72%, and 11.28% across the three survey datasets. The ‘GACC’ variable reflects respondents’ perceptions of government actions (GA) to mitigate global climate change (CC). This variable is divided into three categories: ‘too much’, ‘about the right amount’, and ‘too little’ (with ‘too little’ serving as the reference category). Table 1 indicates that in each sample, more than half of the respondents feel that government efforts to combat global climate change are insufficient.
@@
We also incorporate a broad set of demographic variables as covariates in our regression models. These variables (except marital status) are common across all surveys and include: income, age, gender, education, marital status (not available in the April 2021 survey), metropolitan area, region of residence, birth country, racial background, and party affiliation. We now examine the construction of these variables in more detail.
@@
Pew surveys record income as belonging to one of nine categories: $0-30k,30k-40k,40k-50k,50k-60k,60k-70k,70k-80k,80k-90k,90k-100k$, and $>100k$, where $k$ denotes a thousand dollars. To simplify, we use the mid-point of each income range and assign an imputed value of \$20,000 for the first category and \$110,000 for the last category. Age is recorded into four categories (18-29, 30-49, 50-64, and $\geq 65$ years), but for convenience we recode it to a binary indicator taking the value 1 if the respondent is below 50, and 0 otherwise. As shown in Table 1, 46.36% of respondents in the first dataset belong to the younger age group, with this percentage being approximately 10% lower in the following two datasets. Gender is represented by a binary indicator ‘Female’, where 1 indicates female and 0 indicates male. We see that more than 50% respondents across all three datasets are female. Education is classified into three categories: ‘Graduate & Above’, ‘Some College’, and ‘High School (HS) and Below’; the latter is used as the reference category. As shown in Table 1, the majority of respondents are graduates or have higher education, followed by those with some college education. Therefore, respondents across the three samples are generally well-educated.
@@
Information on marital status is available only in the May 2022 and May--June 2023 surveys. We create an indicator variable ‘Married’ that takes value 1 if the respondent is married or living with a partner, and 0 if they are divorced, separated, widowed, or never been married. As per Table 1, more than three-quarters of respondents are married or living with a partner. Two variables reflect a respondent’s residential background. The ‘Metropolitan’ variable is an indicator that equals 1 if the respondent lives in a metropolitan area, and 0 otherwise. Across the three datasets, more than four-fifths of the respondents are located in metropolitan areas. Additionally, we include regional indicators for ‘Northeast’, ‘Midwest’, ‘South’, and ‘West’; with Northeast serving as the reference category in our regression models. As shown in Table 1, the largest proportion of respondents (about 40%) reside in the South, followed by approximately 22% in both the Midwest and West. The smallest fraction, about 15%, lives in the Northeast.
@@
4 Empirical Results and Discussion
In this section, we report the results from the Bayesian estimation of ordinal probit and ordinal quantile models. We also report the marginal likelihood which may be used for model comparison. Morevoer, for each model, we compute the covariate effects for all or some of the significant variables and discuss our findings in the context of existing literature.
@@ 4.1 Results from Ordinal Probit Models
We begin by presenting the results from ordinal probit models estimated separately for each survey data. Specifically, Table 2 presents the posterior means and standard deviations of the parameters along with the logarithm of marginal likelihood (Chib and Jeliazkov, 2001). The estimates are obtained from 10,000 MCMC iterations (after a burn-in of 2,500 iterations) and utilizes the following prior distributions: $\beta\sim N_{k}(0,I)$ and $\delta\sim N_{J-2}(0,0.25\ast I)$, where $I$ denotes an identity covariance matrix. The priors are relatively diffuse to allow the data speak for itself. In ordinal models, the sign of coefficients indicates the direction of covariate effects for the first and last response categories. For example, in Table 2, the coefficient for “EV Info” is positive (0.43), which implies that it increases (decreases) the probability of last (first) category “Very likely” (“Not at all likely”) to purchase an EV. However, these coefficients do not represent the covariate effects nor can we state anything a-priori about the covariate effects on intermediate outcomes (i.e., “Not too likely” and “Somewhat likely”). Therefore, Table 3 provides the covariate effects for all variables which are statistically different from zero. Note, our measure of covariate effects account for uncertainty in parameters and observations to avoid biases (Jeliazkov and Vossmeyer, 2018).
@@
The combined results from the 2021, 2022, and 2023 surveys reveal consistent patterns in factors associated with EV purchase intent. As shown in Table 3, the covariate effects corresponding to the 2021 survey show that respondents who have read or heard extensively about EVs are 10.2 (2.2) percentage point more likely to express strong interest (some interest) in purchasing EVs as compared to those with limited exposure to EV-related information. These gains in probability comes from a reduction of 8.5 and 3.9 percentage points from the first two opinion categories. Our findings align with Zhao et al. (2022) and Scherrer (2023), who demonstrate that greater exposure to information about EVs—whether through peer effects or media—positively influences attitudes toward EVs, and with Krause et al. (2013), who underscore the crucial role of accurate information in shaping consumer perceptions.
@@
Going forward, we discuss the covariate effects only on the first and last categories as they show the largest changes in probability. For the 2021 survey, we also observe, the belief that EVs are environmentally better than gas-powered vehicles is associated with a 15.5 percentage point increase in strong interest and a 19.5 percentage point decrease in no interest, underscoring the importance of environmental perceptions in shaping purchase intent. This finding aligns with Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025) who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs  [^4] Relatedly, Buhmann and Criado (2023) find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
@@
Table 2: Results from ordinal probit models: posterior mean (Mean) and standard deviation (Std) of the parameters, and logarithm of marginal likelihood, from April 2021; May 2022; and May–June 2023 survey data.
@@
Intercept $-0.28$ $0.07$ $0.36$ $0.09$ $-0.05$ $0.09$
EV Info $0.43$ $0.02$.absent... . ..absent... . ..absent... . ..absent... .
Env Better $0.79$ $0.03$.absent... . ..absent... . ..absent... . ..absent... .
EV Infra ..absent... . ..absent... . ..absent... . ..absent... . $0.82$ $0.04$
@@
US Born $-0.21$ $0.04$ $-0.33$ $0.04$ $-0.22$ $0.04$
Black $-0.28$ $0.04$ $-0.14$ $0.06$ $-0.18$ $0.05$
Other Races $0.05$ $0.03$ $0.01$ $0.04$ $-0.03$ $0.04$

### v2

@@ 2 Data
The study draws on survey data from three opinion polls carried out by the Pew Research Center over a three-year span. The first survey was conducted during April 20-21, 2021 on a sample of 13,749 adults, the second during May 2-8, 2022 on a sample of 10,282 adults, and the third during May 30-June 4, 2023 on a sample of 10,329 adults. After removing all missing observations and cleaning the data, we have a sample of 11,052, 7,173, and 7,201 observations available for analysis [^5] All survey data and additional information on the surveys used in this paper are available for download from the Pew Research Center dataset portal (and the corresponding author’s webpage).. Table 1 shows a descriptive summary of the independent and dependent variables. Specifically, it presents the mean and standard deviation of the continuous variables, and count and percentage for the categorical variables.
The data utilized offers at least three distinct advantages over existing studies that use survey data to gauge EV purchase intent. First, they are nationally representative and span multiple years. As a result, the estimates we generate are more likely to be robust to cross-sectional heterogeneity. Second, the availability of ordinal responses to likelihood of purchasing EV enables us to model public preference regarding EV purchase in much greater detail as compared to studies that use binary responses, such as (Buhmann and Criado, 2023). Third, the first and second surveys correspond to the Covid-19 period, while the third survey was conducted in the post-pandemic phase. Therefore, if there was an increase in interest for sustainable products like electric vehicles during the pandemic, this should be evident in the data summary, with a higher percentage of respondents expressing a preference for buying electric vehicles during the pandemic phase.
@@
EV Info $\circ$ Read or heard a lot 4170 37.73 .... .... .... ....
$\circ$ Little or nothing at all 6882 62.27 .... .... .... ....
Env Better $\circ$ EVs are environmentally ‘better’ than gas-powered vehicles 7944 71.88 .... .... .... ....
$\circ$ About the same or worse 3108 28.12 .... .... .... ....
EV Infra $\circ$ US will build charging stations and infrastructure: extremely or very confident .... .... .... .... 1263 17.54
$\circ$ Somewhat confident, not too confident, or not at all confident .... .... .... .... 5938 82.46
@@
Married Married or living with a partner .... .... 5692 79.35 5640 78.32
@@
We begin by discussing the four independent variables that are of primary interest in this article. Two such variables, ‘EV Info’ and ‘Env Better’, are specific to the April 2021 survey. The variable ‘EV Info’ measures how much a respondent has read or heard about EVs. It is coded as 1 if the respondent has read or heard a lot about EVs, and 0 if they have read or heard little or nothing. As shown in Table 1, about 38% of respondents have read or heard a lot about EVs. The variable ‘Env Better’ reflects perception about the environmental impact of EVs as compared to gas-powered vehicles. This indicator variable takes the value 1 if the respondent believes that EVs are better for the environment than gas-powered vehicles, and 0 otherwise (worse for the environment or about the same). According to Table 1, approximately 72% of respondents view EVs as being environmentally superior to gas-powered vehicles.
@@
The third variable of interest ‘EV Infra’ is only available in the May-June 2023 survey. This indicator variable is coded as 1 if the respondent is extremely or very confident that the US will build the necessary charging stations and infrastructure to support large number of EVs on the roads, and 0 otherwise (somewhat confident, not too confident, or not at all confident). As seen in Table 1, about 82% of respondents express a negative perception towards the government for creating the infrastructure needed to support the adoption of EVs. The fourth variable of interest ‘GACC’ reflects respondents’ perceptions of government actions (GA) to mitigate global climate change (CC) and is available in all three surveys. This variable is divided into three categories: ‘too much’, ‘about the right amount’, and ‘too little’ (with ‘too little’ serving as the reference category). Table 1 indicates that in each sample, more than half of the respondents feel that government efforts to combat global climate change are insufficient.
@@
Another variable, which is an important predictor for EV purchase, is EV Owner, indicating whether the respondent currently owns an electric or hybrid vehicle and taking the value 1 if true and 0 otherwise. As shown in Table 1, the percentages of respondents who are current owners of an electric or hybrid vehicle are 9.40%, 10.72%, and 11.28% across the three survey datasets. We also incorporate a broad set of demographic variables as covariates in our regression models. These variables (except marital status) are common across all surveys and include: income, age, gender, education, marital status (not available in the April 2021 survey), metropolitan area, region of residence, birth country, racial background, and party affiliation. We now examine the construction of these variables in more detail.
@@
Pew surveys record income as belonging to one of nine categories: 030k,30k40k,40k50k,50k60k,60k70k,70k80k,80k90k,90k100k0-30k,30k-40k,40k-50k,50k-60k,60k-70k,70k-80k,80k-90k,90k-100k, and >100k>100k, where $k$ denotes a thousand dollars. To simplify, we use the mid-point of each income range and assign an imputed value of \$20,000 for the first category and \$110,000 for the last category. Age is recorded into four categories (18-29, 30-49, 50-64, and $\geq 65$ years), but for convenience we recode it to a binary indicator taking the value 1 if the respondent is below 50, and 0 otherwise. As shown in Table 1, 46.36% of respondents in the first dataset belong to the younger age group, with this percentage being approximately 10% lower in the following two datasets. Gender is represented by a binary indicator ‘Female’, where 1 indicates female and 0 indicates male. We see that more than 50% respondents across all three datasets are female. Education is classified into three categories: ‘Graduate & Above’, ‘Some College’, and ‘High School (HS) and Below’; the latter is used as the reference category. As shown in Table 1, the majority of respondents are graduates or have higher education, followed by those with some college education. Therefore, respondents across the three samples are generally well-educated.
@@
Information on marital status is available only in the May 2022 and May-June 2023 surveys. We create an indicator variable ‘Married’ that takes value 1 if the respondent is married or living with a partner, and 0 if they are divorced, separated, widowed, or never been married. As per Table 1, more than three-quarters of respondents are married or living with a partner. Two variables reflect a respondent’s residential background. The ‘Metropolitan’ variable is an indicator that equals 1 if the respondent lives in a metropolitan area, and 0 otherwise. Across the three datasets, more than four-fifths of the respondents are located in metropolitan areas. Additionally, we include regional indicators for ‘Northeast’, ‘Midwest’, ‘South’, and ‘West’; with Northeast serving as the reference category in our regression models. As shown in Table 1, the largest proportion of respondents (about 40%) reside in the South, followed by approximately 22% in both the Midwest and West. The smallest fraction, about 15%, lives in the Northeast.
@@
3 The Modeling Framework
3.1 The Ordinal Probit Model
Ordinal data models (Johnson and Albert, 2000), also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates. These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman (2012) and Batham et al. (2023). In a typical setting, the response variable $y$ has several categories and each category is assigned a score (value or number) which is inherently ordered or ranked, without any cardinal interpretation. For example, in our study, responses ($y$) to the question about purchasing an EV are coded as follows: 1 for ‘not at all likely’, 2 for ‘not too likely’, 3 for ‘somewhat likely’, and 4 for ‘very likely’. Here, a score of 2 as compared to 1 implies more inclination to buy an electric vehicle, but we cannot interpret a score of 2 as twice the inclination compared to a score of 1. Similarly, the difference in inclination between the scores 2 and 1 is not the same as that between 4 and 3.
In our study, the unobserved variable $z_{i}$ represents an individual’s intent to purchase EV and is linked to the observed ordinal response $y_{i}$ as follows:
$\gamma_{j-1}<z_{i}\leq\gamma_{j}\;\longleftrightarrow\;\emph{$y_{i}$ = j},\hskip 54.2025pt\forall\;i=1,\cdots,n;\;j=1,\cdots,J,$ (2)
where $\infty=\gamma_{0}<\gamma_{1}\cdots<\gamma_{J-1}<\gamma_{J}=\infty$ are the cut-points (or thresholds) and $J$ denotes the number of categories or outcomes of $y$. This relationship between the latent $z$ (intent to purchase an EV) and observed $y$ (response to the question about purchasing an EV) is illustrated in Figure 2. As shown, $P(y_{i}=1)$ is the area under $f(z_{i})$ to the left of $\gamma_{1}$, $P(y_{i}=2)$ is the area under $f(z_{i})$ between $\gamma_{1}$ and $\gamma_{2}$, and so forth. To specify the location and scale restrictions, we fix $\gamma_{1}=0$ and $var{(\varepsilon)}=1$, respectively (Jeliazkov et al., 2008; Jeliazkov and Rahman, 2012).
[…]
Given the responses $y$ = $(y_{1},\cdots,y_{n})^{\prime}$ and the covariates, the likelihood function for ordinal probit model can be written as follows:
where $\Phi(\cdot)$ is the cumulative distribution function (cdf) of a standard normal distribution and $I(y_{i}=j)$ is an indicator function, which equals 1 if the condition within parenthesis is true and 0 otherwise. But the ordering constraint $\gamma_{1}=0<\gamma_{2}<...<\gamma_{J-1}$ is difficult to satisfy during the MCMC sampling. So, following Albert and Chib (2001), we apply the transformation $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $2\leq j\leq J-1$, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
To complete the Bayesian framework, we utilize the Bayes theorem to combine the complete data likelihood $f(y,z|\beta,\delta)$ with the prior distributions: $\beta\sim N\left(\beta_{0},B_{0}\right)$ and $\delta\sim N\left(d_{0},D_{0}\right)$. The resulting complete data posterior can be written as,
where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, with the index $j$ determined by $y_{i}$, $f(z|\beta)=f_{N}(z|X\beta,I_{n})$, and $\pi(\beta)$ and $\pi(\delta)$ are the normal priors. Given the complete data posterior (4), the conditional posteriors can be derived and the model estimated using MCMC sampling as outlined in Algorithm 1.
Post estimation, an important construct is to compute the covariate effects (or marginal effects) on the outcome probabilities, since, unlike linear regression, the coefficients do not give the covariate effects. Let the $l$-th covariate, $x_{i,l}$}, take two distinct values, $a$ and $b$, denoted as $x_{i,l}^{a}$} and $x_{i,l}^{b}$}, respectively. We partition the covariate and parameter vectors as follows: $x_{i}^{a}=(x_{i,l}^{a},x_{i,-l})$, $x_{i}^{b}=(x_{i,l}^{b},x_{i,-l})$, and $\beta=(\beta_{l},\beta_{-l})$, where l-l denotes all covariates (and parameters) except for the $l$-th covariate (and parameter). We evaluate the distribution of the difference $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$} for $1\leq j\leq J$, by marginalizing over $x_{i,-l}$} and $(\beta,\delta)$. This enables us to account for sampling and parameter uncertainty, and avoid nontrivial biases in estimating covariate effects (Jeliazkov and Vossmeyer, 2018).
Algorithm 1: MCMC algorithm for estimating ordinal probit model.
  1. 1.
    Sample $\delta,z|y,\beta$ in one block as follows:
    1. (a)
[…]
      otherwise repeat the current value $\delta$.
    2. (b)
  2. 2.
[…]
To obtain draws from the distribution $\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})$}, we use the method of composition (Jeliazkov et al., 2008). This involves selecting an individual, extracting the corresponding covariate values, drawing $(\beta,\delta)$ from their posterior distributions, and evaluating the difference $\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta,\delta)-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta,\delta)$}, where $\Pr(y_{i}=j|x_{i}^{q},\beta,\delta)=\Phi(\gamma_{j}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l})-\Phi(\gamma_{j-1}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l}),$ for $q=b,a$, and $1\leq j\leq J$. The process is repeated for all individuals and MCMC draws. The average covariate effect (ACE) for outcome $j$ is then computed as the mean of pointwise probability differences:
3.2 The Ordinal Quantile Model
The parameter estimates from the ordinal probit model provide information on the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates. However, researchers are often more interested in examining specific quantiles of the dependent variable, particularly the lower and upper extremes of the distribution, to answer policy-relevant questions. For instance, having confidence that public charging infrastructure will continue to expand—one of our key covariates—may have less effect on individuals with low interest in EVs. However, among those who are moderately interested, increased confidence in future infrastructure may strengthen intent, reassuring them that charging access will not be a barrier and moving them from “somewhat likely” to “very likely” buyers. To capture how the influence of infrastructure confidence and other factors varies across levels of purchase intent, we use ordinal quantile regression.
$\gamma_{p,j-1}$ $<z_{i}\leq\gamma_{p,j}\;\longleftrightarrow\;\emph{$y_{i}$ = j},$ $\qquad\forall\;i=1,\cdots,n;\;j=1,\cdots,J,$
At this stage, three key issues need attention. First, the ordering constraint of cut-points during sampling is addressed using the transformation, $\delta_{p,j}=\ln(\gamma_{p,j}-\gamma_{p,j-1})$ for $2\leq j\leq J-1$. Second, we set $\gamma_{p,1}=0$ to anchor the location and variance is automatically fixed for any quantile to meet the scale restriction. Third, the AL distribution is not convenient for MCMC sampling as it does not yield tractable conditional posteriors. So, following Kozumi and Kobayashi (2011), we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$. The variable $w_{i}$ follows an exponential distribution, $w_{i}\sim\mathcal{E}(1)$, and is independent of $u_{i}$, which follows a normal distribution, $u_{i}\sim N(0,1)$. Based on this formulation, $z_{i}|\beta_{p},w_{i}\sim N(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\ldots,n$; enabling us to leverage the properties of normal distribution for an efficient MCMC algorithm.
By Bayes theorem, we combine the complete data likelihood $f(y,z|\beta_{p},\delta_{p},w)$ with the normal priors: $\beta_{p}\sim N(\beta_{p0},B_{p0})$, $\delta_{p}\sim N(\delta_{p0},D_{p0})$, to arrive at,
Algorithm 2: MCMC algorithm for estimating ordinal quantile model.
  1. 1.
    Sample $\beta_{p}|z,w$ $\sim$ $N(\tilde{\beta}_{p},\tilde{B}_{p})$, where,
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
The conditional posterior distributions are derived from the complete data posterior (i.e., equation 8) and the parameters are sampled as per Algorithm 2. The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman (2023).
After estimating the ordinal quantile models, we use the method of composition to compute the covariate effects, such as in (Rahman and Vossmeyer, 2019) and Bresson et al. (2021). Assuming a setup similar to Section 3.1, our goal is to draw from the distribution $\{\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})\}$. So, we randomly select an individual, extract their covariates, draw a value $(\beta_{p},\delta_{p})$ from their posterior distributions, and finally evaluate $\{\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta_{p},\delta_{p})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta_{p},\delta_{p})\}$, where now for the quantile model: $\Pr(y_{i}=j|x_{i}^{q},\beta_{p},\delta_{p})=F_{AL}(\gamma_{p,j}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})-F_{AL}(\gamma_{p,j-1}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})$, for $q=b,a$ and $1\leq j\leq J$. This process is repeated for all individuals and MCMC draws, and the average covariate effect for outcome $j$ at quantile $p$ is calculated as the mean of difference in pointwise probabilities.
4 Empirical Results
@@ 4.1 Results from Ordinal Probit Models
We begin by presenting the results from ordinal probit models estimated separately for each survey. Table 2 presents the posterior means and standard deviations of the parameters based on 10,000 MCMC iterations after a burn-in of 2,500 iterations. We utilize the priors $\beta\sim N_{k}(0,I)$ and $\delta\sim N_{J-2}(0,0.25\times I)$, where $I$ denotes an identity covariance matrix. These priors are relatively diffuse, allowing the data to primarily drive the posterior inference. Table 2 also reports the logarithm of marginal likelihood, which is the most accepted quantity for Bayesian model comparison (Chib and Jeliazkov, 2001; Jeliazkov et al., 2008). The marginal likelihood (also known as model evidence) is the probability of observed data, obtained by integrating the likelihood function with respect to prior distribution. It accounts for both model fit and complexity, with higher values indicating a better model fit. The use of marginal likelihood contrasts with those commonly used in the classical framework, such as the pseudo-$R^{2}$, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) (Lin and Tan, 2017; Kumar et al., 2025).
In ordinal models, coefficient signs indicate the direction of covariate effects for the first and last response categories. For example, in Table 2, ‘EV Info’ has a positive coefficient (0.43), so it increases (decreases) the probability of last (first) category “Very likely” (“Not at all likely”). However, these coefficients do not represent the covariate effects nor can we state anything a-priori about the effects on intermediate outcomes (“Not too likely” and “Somewhat likely”). Therefore, Table 3 provides the covariate effects for all variables whose parameters are statistically different from zero at 95% probability level. [^6] As noted earlier, some variables of interest are available in only one year. For variables measured in all three survey years, however, the estimated effects are similar across years, suggesting that the influences on EV purchase intent are relatively stable over time. This stability supports the credibility of estimates for variables observed in a single year.
@@
The combined results from the three surveys reveal consistent patterns in factors associated with EV purchase intent. As shown in Table 3, the covariate effects from the 2021 survey show that respondents who have read or heard extensively about EVs are 10.2 (2.2) percentage point more likely to express strong interest (some interest) in purchasing EVs, compared to those with limited exposure. These gains in probability comes from a reduction of 8.5 and 3.9 percentage points from the first two opinion categories. Our findings complements Zhao et al. (2022) and Scherrer (2023), who demonstrate that greater exposure to EV information through peers or media positively influences attitudes toward EVs, and with Krause et al. (2013), who underscore the role of accurate information in shaping perceptions.
@@
Going forward, we discuss the covariate effects only on the first and last categories as they show the largest changes in probability. For the 2021 survey, the belief that EVs are environmentally better than gas-powered vehicles is associated with a 15.5 percentage point increase in strong interest and a 19.5 percentage point decrease in no interest, underscoring the importance of environmental perceptions in shaping purchase intent. This finding complements Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025) who show that environmentally sensitive respondents tend to have a more favorable opinion of EVs. Here the effect reflects perceptions of EVs’ benefits over gasoline vehicles rather than general pro-environmental attitudes [^7] Relatedly, Buhmann and Criado (2023) find that reputation-conscious consumers prefer EVs only when their prices exceed those of internal combustion engine vehicles, suggesting that, for these consumers, reputation takes precedence over environmental concerns..
@@
Table 2: Results from ordinal probit models: posterior mean (Mean) and standard deviation (Std) of the parameters, and logarithm of marginal likelihood ($\ln$ ML), from April 2021; May 2022; and May–June 2023 survey data.
@@
Intercept 0.28-0.28 $0.07$ $0.36$ $0.09$ 0.05-0.05 $0.09$
EV Info $0.43$ $0.02$ .... .... .... ....
Env Better $0.79$ $0.03$ .... .... .... ....
EV Infra .... .... .... .... $0.82$ $0.04$
GACC (TM) 0.79-0.79 $0.04$ 1.13-1.13 $0.04$ 1.07-1.07 $0.04$
GACC (RA) 0.52-0.52 $0.03$ 0.63-0.63 $0.03$ 0.49-0.49 $0.04$
@@
US Born 0.21-0.21 $0.04$ 0.33-0.33 $0.04$ 0.22-0.22 $0.04$
Black 0.28-0.28 $0.04$ 0.14-0.14 $0.06$ 0.18-0.18 $0.05$
Other Races $0.05$ $0.03$ $0.01$ $0.04$ 0.03-0.03 $0.04$